OpenStudy (anonymous):

how do you solve csc^2(2x)=4/3

OpenStudy (btaylor):

csc x = 1/sin(x), so change csc to sin by: \[\csc^2 (2x) = \frac{4}{3} \rightarrow \frac{1}{\sin^2 (2x)}=\frac{4}{3} \rightarrow \sin^2 (2x)=\frac{3}{4}\] Can you figure it out from here?

OpenStudy (anonymous):

Square root then inverse sine?